# Bose-Einstein condensation in exotic lattice geometries

### v1

@@ Abstract
Modern quantum engineering techniques allow for synthesizing quantum systems in exotic lattice geometries, from self-similar fractal networks to negatively curved hyperbolic graphs. We demonstrate that these structures profoundly reshape Bose–Einstein condensation. Fractal lattices dramatically lower the condensation temperature, while hyperbolic lattices cause it to increase as the system grows – a behavior not seen in ordinary two-dimensional arrays, where the condensation temperature vanishes in the large-size limit. The underlying geometry also controls condensate fluctuations, enhancing them on fractal networks but suppressing them on hyperbolic graphs compared with regular one-dimensional or two-dimensional lattices. When strong repulsive interactions are included, the gas enters a Mott insulating state. A multi-site Gutzwiller approach finds a smooth interpolation between the characteristic insulating lobes of one-dimensional and two-dimensional systems. Re-entrant Mott transitions are seen within a first-order resummed hopping expansion. Our findings establish lattice geometry as a powerful tuning knob for quantum phase phenomena and pave the way for experimental exploration in photonic waveguide arrays and Rydberg-atom tweezer arrays.
@@ I Introduction
Figure 1: Illustrations of lattice geometries used in this work: (a) Sierpiński gasket with dimension $d\approx 1.585$, (b) triangular lattice with $d=2$, (c) Sierpiński carpet with $d\approx 1.893$, (d) square lattice with $d=2$, (e) Sierpiński tetrahedron with $d=2$, and (f) tetrahedral lattice with $d=3$, as well as hyperbolic lattices with: $\{p,q\}=\{3,7\}$, with $V=96$ sites (g), $V=4264$ (i) and $\{p,q\}=\{7,3\}$, with $V=112$ (h), and $V=3481$ (j) sites. Lattices on Fig 1(i,j) have a vertex as their center, while lattices on Fig 1(g,h) are constructed around a central $p$-polygon.
@@
In 1995, the experimental discovery of Bose-Einstein condensation in a gas of Sodium [1] or Rubidium atoms [2] has been a ground-breaking achievement: It has confirmed the theoretical prediction made by Bose [3] and Einstein [4] 70 years earlier, and at the same time, it has opened a new research field that uses cold atoms in order to scientifically explore and technologically exploit quantum phenomena. The phenomenon of Bose-Einstein condensation has not remained limited to atomic gases: Not only does the superfluid phase of a 4He liquid, which has been known since 1937 already, carry many characteristic signatures of a Bose-Einstein condensate (BEC), but bosonic quasiparticle can also be condensed into a BEC. Quasiparticle BECs have been realized a few years after the first atomic BECs with magnons [5], excitons [6], or exciton-polaritons [7, 8]. Finally, also the condensation of elementary bosons is possible: In 2010, the first BEC of photons has been achieved [9]. The macroscopic occupation of the single-particle ground state forms the basis of Bose-Einstein condensation. It goes hand in hand with a series of fascinating properties [10]: The condensate exhibits long-range phase coherence, which in the presence of weak interactions turns into superfluid behavior. This can be evidenced, for instance, through the presence of vortices [11, 12]. Thermal effects and/or interactions deplete the condensate, and peculiar fluctuations of the ground state occupation have been extensively studied theoretically [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28], and experimentally [29, 30, 31, 32, 33, 34, 35, 36, 37, 38], for recent review see [39]. The existence of a condensate also depends strongly on geometric properties of the system. For instance, a uniform gas of massive bosons cannot condense at non-zero temperature in less than three spatial dimensions, but also 1D or 2D systems can show condensation in the presence of trapping potentials or in finite systems [10]. The present manuscript revisits the phenomenon of Bose-Einstein condensation from the point of view of exotic geometries, in particular in fractal lattices, characterized by (possibly) non-integer fractal dimensions, or in hyperbolic lattices, characterized by negative curvature.
@@
The motivation behind this work stems from the recent progress in quantum engineering techniques which has provided us with various synthetic quantum systems in such exotic spaces [40]. This includes photonic fractal lattices [41, 42], synthetic electronic lattices with fractal structure [43, 44], fractal lattices of cold atoms in optical tweezers [45], hyperbolic lattices realized with superconducting qubits [46]. Many interesting aspects of quantum behavior in unconventional geometries have already been revealed: For instance, in quasiperiodic or fractal structures, the absence of Bloch theorem can give rise to localized or critical eigenstates [47], as has theoretically been known since the early 1980s for 1D quasi-crystals [48, 49] or Sierpiński fractals [50, 51, 52]. Transport behavior and localization phenomena in fractal lattices have been theoretically studied both on the classical level [53, 54, 55, 56, 57], and in the quantum regime [58, 59, 60, 61, 62], including also topological transport behavior [63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77]. Beyond the single-particle picture, also the BCS pairing behavior of fermions in fractal lattices [78] and their Mott transition [79, 80] has been studied. For weakly interacting bosons, the loop current behavior of mini-condensates in a Sierpiński gasket has been analyzed [81].
@@
Similarly, there has been growing interest in hyperbolic lattices as a platform for studying quantum phenomena in non-Euclidean geometries. Recent research has explored topological states [82, 83, 84, 85, 86], the Bose-Hubbard model [87], quantum phase transitions [88] and Bose-Einstein condensation [89] in hyperbolic spaces, as well as other curved geometries [90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105]. Studies have also investigated the density of states in hyperbolic tight-binding models [106] and developed a framework for the crystallography of hyperbolic lattices [107, 108].
@@
In this work, we focus on fractal and hyperbolic lattices, as examples of exotic geometries that extend the more commonly investigated Euclidean two-dimensional lattice types. The geometries will be introduced in detail in Sec. II. We then consider both quantum-statistical effects of non-interacting bosons, and interaction effects in a Bose-Hubbard model. The condensation behavior seen in these geometries is compared to the known behavior in regular lattices [109], highlighting the important effect of geometry on captivating many-body quantum phenomena.
@@
Sec. III concentrates on a non-interacting boson gas at finite temperature. We compare various properties of the Bose-Einstein condensate in regular, fractal and hyperbolic geometries, including the scaling of condensate fraction and condensate fluctuations, off-diagonal long-range order, and dependence of critical temperature on system size. We find that the condensate fraction in fractal lattices drops to zero at absolute temperatures much lower than in the corresponding regular lattices with the same number of sites. The finite size scaling of the critical temperature indicates that condensation in fractal lattices occurs only in finite systems, as is also the case in regular 2D lattices (cf. [110, 111]). Interestingly, this observation also holds for the Sierpiński tetrahedron lattice, which is embedded in three dimensions. The situation is strikingly different in hyperbolic lattices – we find that, despite being embedded in 2D, the condensate fraction scales with temperature similarly to the regular 3D lattice. Importantly, the critical temperature does not decrease monotonically with the system size, which suggests a non-zero $T_{c}$ in the thermodynamic limit. We also analyze the fluctuations above the condensate, which are found to be enhanced in the fractal geometries, but strongly suppressed in the hyperbolic lattice, as compared to regular lattices in 2D or 3D.
@@
In Sec. IV, we investigate the properties of interacting Bose gas at zero temperature. We consider the Bose-Hubbard model in graphs corresponding to lattices with different geometries (cf. [112, 113, 80]). We investigate the Mott insulator (MI) to superfluid (SF) phase transition as a function of the hopping amplitude $J$ and chemical potential $\mu$ [114], using both a Green function method [115, 116, 117] and a multi-site Gutzwiller approach [118, 119, 120, 121]. In contrast to the single-site Gutzwiller approach, which simplifies the many-body problem into a self-consistent local problem, the multi-site approach is able to capture also properties of the lattice geometry beyond the average coordination number $z$. Specifically, the Gutzwiller calculation yields Mott lobes for the Sierpiński triangle that nicely illustrates the intermediate dimensionality of the system: the lobe exhibits a kink at the tip, as also found in 1D systems [122, 123], while the shoulders of the lobe remain convex, as for the lobes in regular 2D lattices [124]. Based on a resummed hopping expansion, the Green function method predicts a re-entrant SF to Mott transition for the Sierpiński triangle, due to the presence of spectral gaps in the tight-binding band structure. Since the multi-site Gutzwiller method does not show such a re-entrant behavior, we tend to interpret it as an artifact of the hopping expansion.
@@ II Model
We start with a lattice defined in the language of a mathematical graph. A lattice ${\cal L}$ is a connectivity graph ${\cal L}=({\cal V},{\cal E})$ where ${\cal V}=\{1,\dots,V\}$ is a set of enumerated nodes, while =${\cal E}=\{\langle i,j\rangle\in{\cal V}\times{\cal V}|i\neq j\}$ is a set of edges on the lattice. The set of edges defines an adjacency matrix of a graph $J_{ij}=J_{ji}$, $i,j=1,\dots,L$, which takes $1$ for any pair $\langle i,j\rangle\in{\cal E}$, and zero otherwise. The tight-binding Hamiltonian on the lattice ${\cal L}$ reads
$\hat{H}_{\cal L}^{0}=-J\sum_{i,j\in{\cal V}}J_{ij}(\hat{b}_{i}^{\dagger}\hat{b}_{j}+\hat{b}_{j}^{\dagger}\hat{b}_{i})-\mu\sum_{i\in{\cal V}}\hat{n}_{i},$ (1)
where $\hat{b}_{i}$, $\hat{b}_{i}^{\dagger}$, $\hat{n}_{i}$ are annihilation, creation, and number operator on site $i$. Here, we have chosen a grand-canonical description, in which the particle number $N=\sum_{i\in{\cal V}}\hat{n}_{i}$ is controlled by a chemical potential $\mu$. The parameter $J$ is the hopping amplitude. In the presence of interactions, the hopping competes with on-site repulsion $U$, and the system is described by the Bose-Hubbard Hamiltonian:
@@
Hyperbolic lattices are constructed from regular tilings of the hyperbolic plane, defined by their Schläfi symbol $\{p,q\}$ [126]. The hyperbolic lattice is constructed with $p$-sided polygons, where the average vertex is connected by $q$ edges. We construct our hyperbolic lattices as graphs with open boundary conditions, starting with either a $p$-sided polygon or a single vertex with $q$ $p$-sided polygons around it, see Fig 1. Larger lattices are constructed by adding a layer of polygons around the smaller lattice. Hyperbolic lattices that satisfy the equation $(p-2)(q-2)>4$, such as the $\{p,q\}=\{7,3\}$ and $\{p,q\}=\{3,7\}$ lattices, are characterized by a constant negative curvature [127].
@@ III Non-interacting boson gas
We first analyze the quantum-statistical behavior of non-interacting bosons in different geometries, including regular, fractal, and curved lattices. To this end, let $\epsilon_{k}/J$ denote the eigenvalues of the adjacency matrix $-J_{ij}$. The non-interacting Hamiltonian in diagonal form reads
@@
Let $k=0$ denote the ground state. If the occupation of the ground state, $n_{0}$, becomes macroscopic, the system is considered to be condensed. For a precise definition of such macroscopic ground state occupation, we evaluate the number of excited particles $N_{\rm ex}(\mu=\epsilon_{0})=\sum_{k>0}n_{k}$ for $\mu=\epsilon_{0}$. Since $n_{0}$ (and hence also the total particle number $N=n_{0}+N_{\rm ex}$ diverges for $\mu=\epsilon_{0}$, in practice $\mu<\epsilon_{0}$, and $N_{\rm ex}(\mu=\epsilon_{0})\equiv N_{\rm ex}^{\rm max}$ is an upper bound for the number of excited particles in the system at the given temperature, $N_{\rm ex}(\mu)<N_{\rm ex}^{\rm max}$. For some interval $\mu_{\rm c}<\mu<\epsilon_{0}$, the number of ground state particles $n_{0}(\mu)$ will exceed the maximum number of excited particles $N_{\rm ex}^{\rm max}$, and the system is considered to be condensed. To the critical chemical potential $\mu_{\rm c}$ corresponds a critical total number of particles $N_{\rm c}=2N_{\rm ex}^{\rm max}$ which depends on the chosen temperature. By inverting this function $N_{\rm c}(T)$, we obtain the critical temperature $T_{\rm c}$ below which condensation sets in for a given particle number $N_{\rm c}(T_{\rm c})$.
@@ III.1 Condensate fraction
Figure 2: Condensate fraction in regular and hyperbolic lattice geometries. We plot the condensate fraction as a function of the temperature in units of the critical temperature $T_{c}$ for (a) a three dimensional simple cubic lattice with $V=15^{3}$ sites, and (b) a two dimensional square lattice with $V=60^{2}$ sites, both with open boundary conditions, as well as a $\{p,q\}=\{3,7\}$ hyperbolic lattice with $V=6615$ sites (c), and a $\{p,q\}=\{7,3\}$ hyperbolic lattice with $V=5887$ sites (d). The different colored points correspond to different lattice fillings. The solid black line corresponds to the theoretical curve of Eq. (5), for the regular geometries. In the hyperbolic lattices, the condensate fraction behaves similarly to the 3D cubic lattice.
Figure 3: Condensate fraction in fractal lattices vs temperature. Panel (a) shows, on a linear scale, the condensate fraction in a Sierpiński triangle lattice ($V=3282$ sites) in units of the critical temperature. The condensate fraction drops linearly at small temperatures with a heavy tail above $T_{c}$, regardless of the filling. In panels (b-d), the behavior in different fractal lattices is compared to their non-fractal counterparts with a similar number of sites: (b) Sierpiński triangle lattice ($V=3282$ sites) and standard triangular lattice ($V=3321$ sites), (c) Sierpiński carpet ($V=4096$ sites) and a square lattice ($V=4096$ sites), (d) Sierpiński tetrahedron lattice ($V=2050$ sites) to a regular tetrahedral lattice ($V=2024$ sites). In all plots, unit filling is chosen, and the condensate fraction is plotted vs a logarithmic temperature scale in units of the hopping constant $J$. The vertical lines mark the critical temperatures of the fractal and reference lattices at unit filling.
For any temperature, the total number of particles $N$ shall now be fixed (via $\mu$) according to a desired filling $N/V$ of the lattice, where $V$ is the number of sites in the finite lattice. Acccording to the mentioned procedure, we then define $T_{\rm c}$, and evaluate the condensate fraction $n_{0}/N$ as a function of normalized temperature $T/T_{\rm c}$. Within the inverval $0\leq T\leq T_{\rm c}$, The behavior of the condensate fraction can be fit to a function
@@
As shown in Fig. 2(a), a similar behavior characterizes also the ideal gas in a 3D cubic lattice, although at small fillings a slightly slower decay of condensate fraction is observed due to the finite size effects, whereas at larger filling lattice effects cause a slightly faster decay. In a 2D square lattice, the condensate fraction decays much faster (with $\alpha\approx 1$) for any filling, as shown in Fig. 2(b). Remarkably, by changing from an Euclidean plane to a hyperbolic one, the behavior of the condensate fraction becomes similar to the one of the cubic lattice, see Fig. 2(c-d).
@@
In contrast, the condensate behavior in fractal lattices is barely modeled by Eq. (5), see Fig. 3(a) for the case of a Sierpiński triangle. In particular, the condensate fraction in the fractal lattice is characterized by a heavy tail, with non-zero $n_{0}$ even for $T\gg T_{\rm c}$. However, it must also be noted that the critical temperatures on fractal lattices are orders of magnitude smaller than the critical temperatures in regular lattices of comparable size. In Fig. 3(b–d) we compare the behavior of the condensate fraction in a Sierpiński triangle and a regular triangular lattice (b), in a Sierpiński carpet and a regular square lattice (c), and in a Sierpiński tetrahedron and a regular tetrahedral lattice (d), with the vertical lines marking the critical temperatures in the different geometries. This comparison illustrates that very distinct temperature scales are relevant for fractal and regular lattices, and the heavy tail in the fractal lattice extends approximately in the temperature range between the critical temperature of the fractal lattice and the critical temperature of the corresponding regular lattice.
@@
Figure 4: Critical temperature $T_{c}$ at a unit filling as a function of the system size, in linear scale on Figure 4(b) and double logarithmic scale on Figures 4(a,c). Figure (a) compares the Sierpiński gasket ($d\approx 1.585$) and Sierpiński tetrahedron ($d=2$) to regular 1D, triangular and tetrahedral lattices. The $T_{c}$ decreases and decays faster in lattices with smaller dimension, and doesn’t decay in the 3D tetrahedral lattice. Figure (b) shows the critical temperature in cubic lattices, which depending on the type of boundary conditions, approaches the limit from below or above. Figure (c) shows the $T_{c}$ in square lattices with open boundary conditions and in Sierpiński carpet lattices.
@@
In Fig. 3(c) we also observe an interesting behavior: The critical temperature in the carpet is several orders of magnitudes lower than the one in the square lattice. This results in a fast drop of the condensate fraction until it slows down when $n_{0}$ reaches $0.25$. The next drop occurs at temperatures close to the $T_{c}$ of the standard geometry. This behavior can be explained by examining the spectrum of the carpet lattice: The three first excited states lie very close to the ground state (within $4\cdot 10^{-6}J$) while the next excited state is more than $2\cdot 10^{-3}J$ away from the ground state. Therefore, a condensate of the true ground state depletes quickly, and fragments into the macroscopic occupation of the lowest four eigenstates, before the occupation spills towards higher branches of the spectrum. This example nicely illustrates how the fractal nature of the lattice can give rise to fractal structures in the energy spectrum, which then also manifest in the condensation behavior.

### v2

@@ Abstract
Modern quantum engineering techniques allow for synthesizing quantum systems in exotic lattice geometries, from self-similar fractal networks to negatively curved hyperbolic graphs. We demonstrate that these structures profoundly reshape Bose-Einstein condensation. Fractal lattices dramatically lower the condensation temperature and enhance condensation fluctuations. In a Sierpiński carpet, quasi-degeneracies in the tight-binding spectrum fragment the condensate. Hyperbolic lattices, on the other hand, exhibit condensation features similar to regular three-dimensional lattices, despite their embedding in only two dimensions: The critical temperature increases as the system grows, and the temperature-dependence of the condensate fraction follows the same power-law as for cubic lattices. We explain these similarities through the similarity of the densities of state at low energies. When strong repulsive interactions are included, the gas enters a Mott insulating state. Using a multi-site Gutzwiller approach as well as a simple strong-coupling expansion, for the Sierpiński triangle we find a smooth interpolation between the characteristic insulating lobes of one-dimensional and two-dimensional systems. Our findings establish lattice geometry as a powerful tuning knob for quantum phase phenomena and pave the way for experimental exploration in photonic waveguide arrays and Rydberg-atom tweezer arrays.
@@ I Introduction
Figure 1: Illustrations of lattice geometries used in this work: (a) Sierpiński gasket with dimension $d\approx 1.585$, (b) triangular lattice with $d=2$, (c) Sierpiński carpet with $d\approx 1.893$, (d) square lattice with $d=2$, (e) Sierpiński tetrahedron with $d=2$, and (f) tetrahedral lattice with $d=3$, as well as hyperbolic lattices with: $\{p,q\}=\{3,7\}$, with $V=96$ sites (g), $V=4264$ (i) and $\{p,q\}=\{7,3\}$, with $V=112$ (h), and $V=3481$ (j) sites. Lattices in Fig 1(i,j) have a vertex as their center, while lattices in Fig 1(g,h) are constructed around a central $p$-polygon.
@@
In 1995, the experimental discovery of Bose-Einstein condensation in a gas of sodium [1] or rubidium atoms [2] has been a ground-breaking achievement: It has confirmed the theoretical prediction made by Bose [3] and Einstein [4] 70 years earlier, and at the same time, it has opened a new research field that uses cold atoms in order to scientifically explore and technologically exploit quantum phenomena. The phenomenon of Bose-Einstein condensation has not remained limited to atomic gases: Not only does the superfluid phase of a 4He liquid, which has been known since 1937 already, carry many characteristic signatures of a Bose-Einstein condensate (BEC), but bosonic quasiparticles can also be condensed into a BEC. Quasiparticle BECs have been realized a few years after the first atomic BECs with magnons [5], excitons [6], or exciton-polaritons [7, 8]. Finally, also the condensation of elementary bosons is possible: In 2010, the first BEC of photons has been achieved [9]. The macroscopic occupation of the single-particle ground state forms the basis of Bose-Einstein condensation. It goes hand in hand with a series of fascinating properties [10]: The condensate exhibits long-range phase coherence, which in the presence of weak interactions turns into superfluid behavior. This can be evidenced, for instance, through the presence of vortices [11, 12]. Thermal effects and/or interactions deplete the condensate, and peculiar fluctuations of the ground state occupation have been extensively studied theoretically [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28], and experimentally [29, 30, 31, 32, 33, 34, 35, 36, 37, 38], for recent review, see [39]. The existence of a condensate also depends strongly on geometric properties of the system. For instance, a uniform gas of massive bosons cannot condense at non-zero temperature in less than three spatial dimensions, but also 1D or 2D systems can show condensation in the presence of trapping potentials or in finite systems [10]. The present manuscript revisits the phenomenon of Bose-Einstein condensation from the point of view of exotic geometries, in particular in fractal lattices, characterized by (possibly) non-integer fractal dimensions, or in hyperbolic lattices, characterized by negative curvature.
@@
The motivation behind this work stems from the recent progress in quantum engineering techniques which has provided us with various synthetic quantum systems in such exotic spaces [40]. This includes photonic fractal lattices [41, 42], synthetic electronic lattices with fractal structure [43, 44], fractal lattices of cold atoms in optical tweezers [45], hyperbolic lattices realized with superconducting qubits [46]. Many interesting aspects of quantum behavior in unconventional geometries have already been revealed: For instance, in quasiperiodic or fractal structures, the absence of Bloch theorem can give rise to localized or critical eigenstates [47], as has theoretically been known since the early 1980s for 1D quasi-crystals [48, 49] or Sierpiński fractals [50, 51, 52, 53]. Transport behavior and localization phenomena in fractal lattices have been theoretically studied both on the classical level [54, 55, 56, 57, 58], and in the quantum regime [59, 60, 61, 62, 63], including also topological transport behavior [64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79]. Beyond the single-particle picture, also the BCS pairing behavior of fermions in fractal lattices [80] and their Mott transition [81, 82] has been studied. For weakly interacting bosons, the loop current behavior of mini-condensates in a Sierpiński gasket has been analyzed [83].
@@
Similarly, there has been growing interest in hyperbolic lattices as a platform for studying quantum phenomena in non-Euclidean geometries. Recent research has explored topological states [84, 85, 86, 87, 88], fermionic systems [89, 90, 91], quantum phase transitions [92] and Bose-Einstein condensation [93] in hyperbolic spaces, as well as other curved geometries [94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109]. Studies have also investigated the density of states in hyperbolic tight-binding models [110] and developed a framework for the crystallography of hyperbolic lattices [111, 112].
@@
In this work, we focus on fractal and hyperbolic lattices, as examples of exotic geometries that extend the more commonly investigated Euclidean two-dimensional lattice types. The geometries will be introduced in detail in Sec. II. We then consider both quantum-statistical effects of non-interacting bosons, and interaction effects in a Bose-Hubbard model. The condensation behavior seen in these geometries is compared to the known behavior in regular lattices [113], highlighting the important effect of geometry on captivating many-body quantum phenomena. It is worth noting that the influence of lattice inhomogeneity on BEC [114, 115, 116] and quantum properties in the Bose-Hubbard model [117, 118, 119] has been extensively studied.
@@
Sec. III concentrates on a non-interacting boson gas at finite temperature. We compare various properties of the Bose-Einstein condensate in regular, fractal and hyperbolic geometries, including the scaling of condensate fraction and condensate fluctuations, off-diagonal long-range order, and dependence of critical temperature on system size. We find that the condensate fraction in fractal lattices drops to zero at absolute temperatures much lower than in the corresponding regular lattices with the same number of sites. The finite size scaling of the critical temperature indicates that condensation in fractal lattices occurs only in finite systems, as is also the case in regular 2D lattices (cf. [120, 121]). Interestingly, this observation also holds for the Sierpiński tetrahedron lattice, which is embedded in three dimensions. The situation is strikingly different in hyperbolic lattices – we find that, despite being embedded in 2D, the condensate fraction scales with temperature similarly to the regular 3D lattice. Importantly, the critical temperature does not decrease monotonically with the system size, which suggests a non-zero $T_{\rm c}$ in the thermodynamic limit. We also analyze the fluctuations above the condensate, which are found to be enhanced in the fractal geometries, but strongly suppressed in the hyperbolic lattice, as compared to regular lattices in 2D or 3D.
@@
In Sec. IV, we investigate the properties of interacting Bose gas at zero temperature. We consider the Bose-Hubbard model in graphs corresponding to lattices with different geometries (cf. [122, 123, 82]). We investigate the Mott insulator (MI) to superfluid (SF) phase transition as a function of the hopping amplitude $J$ and chemical potential $\mu$ [124], using both a Green function method [125, 126, 127] and a multi-site Gutzwiller approach [128, 129, 130, 131, 132]. In contrast to the single-site Gutzwiller approach, which simplifies the many-body problem into a self-consistent local problem, the multi-site approach is able to capture also properties of the lattice geometry beyond the average coordination number $z$. Specifically, the Gutzwiller calculation yields Mott lobes for the Sierpiński triangle that nicely illustrates the intermediate dimensionality of the system: the lobe exhibits a kink at the tip, as also found in 1D systems [133, 134], while the shoulders of the lobe remain convex, as for the lobes in regular 2D lattices [135]. Based on a resummed hopping expansion, the Green function method predicts a re-entrant SF to Mott transition for the Sierpiński triangle, due to the presence of spectral gaps in the tight-binding band structure. Since the multi-site Gutzwiller method does not show such a re-entrant behavior, we tend to interpret it as an artifact of the hopping expansion.
@@ II Model
[…]
$\hat{H}_{\cal L}^{0}=-J\sum_{i,j\in{\cal V}}J_{ij}\hat{b}_{i}^{\dagger}\hat{b}_{j}-\mu\sum_{i\in{\cal V}}\hat{n}_{i},$ (1)
where $\hat{b}_{i}$, $\hat{b}_{i}^{\dagger}$, $\hat{n}_{i}$ are annihilation, creation, and number operator acting on site $i$. Here, we have chosen a grand-canonical description, in which the particle number $N=\sum_{i\in{\cal V}}\hat{n}_{i}$ is controlled by a chemical potential $\mu$. The parameter $J$ is the hopping amplitude. In the presence of interactions, the hopping competes with on-site repulsion $U$, and the system is described by the Bose-Hubbard Hamiltonian:
@@
Hyperbolic lattices are constructed from regular tilings of the hyperbolic plane, defined by their Schläfli symbol $\{p,q\}$ [137]. The hyperbolic lattice is constructed with $p$-sided polygons, where the average vertex is connected by $q$ edges. We construct our hyperbolic lattices as graphs with open boundary conditions, starting with either a $p$-sided polygon or a single vertex with $q$ $p$-sided polygons around it, see Fig. 1. Larger lattices are constructed by adding a layer of polygons around the smaller lattice. Hyperbolic lattices that satisfy the equation (p2)(q2)>4(p-2)(q-2)>4, such as the $\{p,q\}=\{7,3\}$ and $\{p,q\}=\{3,7\}$ lattices, are characterized by a constant negative curvature [138].
@@ III Non-interacting boson gas
We first analyze the quantum-statistical behavior of non-interacting bosons in different geometries, including regular, fractal, and curved lattices. To this end, let $\epsilon_{k}/J$ denote the eigenvalues of the adjacency matrix $J_{ij}$. The non-interacting Hamiltonian in diagonal form reads
@@
Let $k=0$ denote the ground state. If the occupation of the ground state, $n_{0}$, becomes macroscopic, the system is considered to be condensed. For a precise definition of such macroscopic ground state occupation, we evaluate the number of excited particles $N_{\rm ex}(\mu=\epsilon_{0})=\sum_{k>0}n_{k}$ for $\mu=\epsilon_{0}$. Since $n_{0}$ (and hence also the total particle number $N=n_{0}+N_{\rm ex}$ diverges for $\mu=\epsilon_{0}$, in practice $\mu<\epsilon_{0}$, and $N_{\rm ex}(\mu=\epsilon_{0})\equiv N_{\rm ex}^{\rm max}$ is an upper bound for the number of excited particles in the system at the given temperature, $N_{\rm ex}(\mu)<N_{\rm ex}^{\rm max}$. For some interval $\mu_{\rm c}<\mu<\epsilon_{0}$, the total number of particles $N(\mu)$ will exceed the maximum number of excited particles $N_{\rm ex}^{\rm max}$ and the system is considered to be condensed, as $n_{0}$ must be macroscopic. To the critical chemical potential $\mu_{\rm c}$ corresponds a critical total number of particles $N_{\rm c}=N_{\rm ex}^{\rm max}$ which depends on the chosen temperature. By inverting this function $N_{\rm c}(T)$, we obtain the critical temperature $T_{\rm c}$ below which condensation sets in for a given particle number $N_{\rm c}(T_{\rm c})$.
@@ III.1 Condensate fraction
Figure 2: Condensate fraction in regular and hyperbolic lattice geometries. We plot the condensate fraction as a function of the temperature in units of the critical temperature $T_{\rm c}$ for (a) a three dimensional simple cubic lattice with $V=20^{3}$ sites, and (b) a two dimensional square lattice with $V=90^{2}$ sites, both with open boundary conditions, as well as a $\{p,q\}=\{3,7\}$ hyperbolic lattice with $V=11173$ sites (c), and a $\{p,q\}=\{7,3\}$ hyperbolic lattice with $V=9136$ sites (d). The different colored points correspond to different lattice fillings. The solid black line corresponds to the theoretical curve of Eq. (5), for the regular geometries. In the hyperbolic lattices, the condensate fraction behaves similarly to the 3D cubic lattice.
Figure 3: Condensate fraction in fractal lattices vs temperature. Panel (a) shows, on a linear scale, the condensate fraction in a Sierpiński triangle lattice ($V=3282$ sites) in units of the critical temperature. The condensate fraction drops linearly at small temperatures with a heavy tail above $T_{\rm c}$, regardless of the filling. In panels (b-d), the behavior in different fractal lattices is compared to their non-fractal counterparts with a similar number of sites: (b) Sierpiński triangle lattice ($V=3282$ sites), a standard triangular lattice ($V=3321$ sites) and a 1D lattice ($V=3282$), (c) Sierpiński carpet ($V=4096$ sites) and square lattice ($V=4096$ sites); for the carpet, both the standard condensate and the fragmented condensate (including the four lowest eigenstates) are shown. Panel (d) compares the Sierpiński tetrahedron lattice ($V=2050$ sites) to a regular tetrahedral lattice ($V=2024$ sites). In all plots, unit filling is chosen, and the condensate fraction is plotted vs a logarithmic temperature scale in units of the hopping constant $J$. The vertical lines mark the critical temperatures of the fractal and reference lattices at unit filling.
For any temperature, the total number of particles $N$ shall now be fixed (via $\mu$) according to a desired filling $N/V$ of the lattice, where $V$ is the number of sites in the finite lattice. According to the mentioned procedure, we then define $T_{\rm c}$, and evaluate the condensate fraction $n_{0}/N$ as a function of normalized temperature $T/T_{\rm c}$. Within the interval $0\leq T\leq T_{\rm c}$, The behavior of the condensate fraction can be fit to a function
@@
As shown in Fig. 2(a), a similar behavior characterizes also the ideal gas in a 3D cubic lattice, although at small fillings a slightly slower decay of condensate fraction is observed due to the finite size effects, whereas at larger filling lattice effects cause a slightly faster decay. In a 2D square lattice, the condensate fraction decays much faster (with $\alpha\approx 1$) for any filling, as shown in Fig. 2(b). Remarkably, by changing from an Euclidean plane to a hyperbolic one, the behavior of the condensate fraction becomes similar to that of the cubic lattice, see Fig. 2(c-d).
@@
Quantitatively, the condensate behavior in fractal lattices is similar to the 2D lattice, as over a broad temperature range the condensate fraction decreases linearly with $T$, see Fig. 3(a) for the case of a Sierpiński triangle. However, a pronounced qualitative difference is the heavy tail of the condensate fraction in fractal lattices, with a non-zero $n_{0}$ even for $T\gg T_{\rm c}$. In this context, it should also be noted that the critical temperatures on fractal lattices are orders of magnitude smaller than the critical temperatures in regular lattices of comparable size. In Fig. 3(b–d) we compare the behavior of the condensate fraction in a Sierpiński triangle and a regular triangular lattice (b), in a Sierpiński carpet and a regular square lattice (c), and in a Sierpiński tetrahedron and a regular tetrahedral lattice (d), with the vertical lines marking the critical temperatures in the different geometries. This comparison illustrates that very distinct temperature scales are relevant for fractal and regular lattices, and the heavy tail in the fractal lattice extends approximately in the temperature range between the critical temperature of the fractal lattice and the critical temperature of the corresponding regular lattice.
@@
Figure 4: Critical temperature $T_{\rm c}$ at a unit filling as a function of the system size, on a double logarithmic scale. The results from a cubic, tetrahedral, triangular, square and 1D lattices are compared to results from fractal geometries: the Sierpiński gasket ($d\approx 1.585$), Sierpiński tetrahedron ($d=2$) and Sierpiński carpet ($d\approx 1.893$). In the Sierpiński carpet, the condensate fragments between the lowest 44 eigenstates, which results in rapid decay of $T_{\rm c}$. One can consider BEC through the occupancy of any of these 44 states, which produces different results (cyan). In general, the $T_{\rm c}$ decreases and decays faster in lattices with smaller dimension, and does not decay in the 3D tetrahedral and cubic lattices.
@@
In Fig. 3(c) we also observe an interesting behavior: The critical temperature in the carpet is several orders of magnitudes lower than the one in the square lattice. This results in a fast drop of the condensate fraction until it slows down when $n_{0}$ reaches $0.25$. The next drop occurs at temperatures close to the $T_{\rm c}$ of the standard geometry. This behavior can be explained by examining the spectrum of the carpet lattice: The three first excited states lie very close to the ground state (within $4\cdot 10^{-6}J$) while the next excited state is more than $2\cdot 10^{-3}J$ away from the ground state. Therefore, a condensate of the true ground state depletes quickly and fragments into the macroscopic occupation of the lowest four eigenstates, before the occupation spills towards higher branches of the spectrum. This example nicely illustrates how the fractal nature of the lattice can give rise to fractal structures in the energy spectrum, which then also manifest in the condensation behavior. One can account for this fragmentation by extending the definition of the condensate fraction to include particles occupying any of the four lowest states. As shown in Fig. 3(c), the obtained temperature dependence of the condensate fraction and the corresponding critical temperature closely resemble those of the 2D square lattice, consistent with the carpet’s Hausdorff dimension $d\approx 1.893$.